Need your help! I am a parent of a 9th grade student who is having a problem with his algebra homework. He is working on "Sysytems of Equations", if you could please give my son some help with the following word problem, this may then help with the rest of his homework.
The difference between two numbers is 16. Three times the larger number is seven times the smaller number. What are the numbers?

Answer by acerX(62)

The first equation is a simple one stating that two numbers "x" and "y" when subtracted will equal 16. Written as:
X - Y = 16
The second one gives us the relationship of the two numbers. Remeber "is" means equals!
3 * X = 7 * Y
3X=7Y Divide both sides by 3
X=7Y/3
Pluging in the value of X in the first equation will give us Y.
7Y/3 - Y = 16 Multiply both sides by 3 to eliminate the fraction.
7Y - 3Y = 48 Combine like terms.
4Y =48 Divide both sides by 4.
Y=12
Now, plug in the Y value in the first equation to find X
X - 12 = 16 Add 12 to both sides
X=28
Hope this helped!